Show that if A_(n×n) is invertible then the inverse is unique


1
Expert's answer
2021-09-03T07:33:10-0400

Since A is a "n\times n" matrix, the linear transformation "T:x\rightarrow Ax" is one-to-one

"\Rightarrow"  linear transformation "T:x\rightarrow Ax" is invertible

"\Rightarrow"  A is invertible.

Suppose that "Ax=b" and "Av=Au=0,v\u2260u."

Then "A(x+v)=A(x+u)=b" , whereby "x+v\u2260x+u."

"\therefore" "Ax=b" for some unique "x"  

"\Rightarrow" either there is no "v" such that "Av=0" or there is a unique "v" such that "Av=0" .

Since "A0=0" , we conclude that there is a unique "v" such that "Av=0" , and thus "T" is one-to-one.

Therefore, A is invertible.